Hi
Mary had 15 coins. There were 8 ten cent coins and the rest were 5c and 20c coins. The total amount was $1.60. How many 5c coins were there.
Thanks

Answer by math_tutor2020(3816)


Mary has 8 coins each valued at 10 cents a piece. That's 8*10 = 80 cents so far. Subtract this from the total $1.60, aka 160 cents, and we get:
160 cents - 80 cents = 80 cents

She has 15 coins total. 8 of which are ten cent coins, so that leaves 15-8 = 7 coins that are either 5 cent or 20 cent coins. These remaining 7 coins account for the remaining 80 cents we calculated in the previous paragraph.

x = number of 5c coins
y = number of 20c coins
x+y = 7
y = 7-x

5x = value from just all the 5c coins only
20y = value from just all the 20c coins only
5x+20y = value from 5c and 20c coins
5x+20y = 80 cents

5x+20y = 80
5x+20(7-x) = 80 .... replace y with (7-x)
5x+140-20x = 80
-15x+140 = 80
-15x = 80-140
-15x = -60
x = -60/(-15)
x = 4

Answer: 4